# ASBench: Image Anomalies Synthesis Benchmark for Anomaly Detection

### v1

@@ Abstract
Anomaly detection plays a pivotal role in manufacturing quality control, yet its application is constrained by limited abnormal samples and high manual annotation costs. While anomaly synthesis offers a promising solution, existing studies predominantly treat anomaly synthesis as an auxiliary component within anomaly detection frameworks, lacking systematic evaluation of anomaly synthesis algorithms. Current research also overlook crucial factors specific to anomaly synthesis, such as decoupling its impact from detection, quantitative analysis of synthetic data and adaptability across different scenarios. To address these limitations, we propose ASBench, the first comprehensive benchmarking framework dedicated to evaluating anomaly synthesis methods. Our framework introduces four critical evaluation dimensions: (i) the generalization performance across different datasets and pipelines (ii) the ratio of synthetic to real data (iii) the correlation between intrinsic metrics of synthesis images and anomaly detection performance metrics , and (iv) strategies for hybrid anomaly synthesis methods. Through extensive experiments, ASBench not only reveals limitations in current anomaly synthesis methods but also provides actionable insights for future research directions in anomaly synthesis.
@@
Anomaly detection, anomalies synthesis, industrial image
@@ 1 Introduction
Anomaly detection has emerged as a critical technology in modern manufacturing quality control and healthcare monitoring, playing an important role in ensuring product reliability and enhancing diagnostic accuracy [5] [29]. Despite its significance, the practical deployment of anomaly detection systems is often constrained by two major challenges: the scarcity of abnormal samples and the prohibitively high costs associated with manual annotation. These limitations significantly impede the performance and scalability of the detection models. To address these issues, anomaly synthesis has gained traction as a promising solution, offering the capability to generate representative abnormal samples. This approach not only alleviates the data scarcity problem but also substantially reduces the reliance on costly manual annotation efforts. Jan and Christian (2022) [7] summerize the anomaly synthesis methods. The significant potential has been demonstrated in various domains, including surface defect detection and medical image analysis. Anomaly synthesis can be effectively integrated with real anomaly data or directly applied to enhance detection systems.
While anomaly synthesis holds transformative potential for advancing anomaly detection capabilities, current anomaly synthesis methods remain largely confinded to auxiliary roles within anomaly detection frameworks. This paradigm neglects rigorous benchmarking and detailed exploration of the methodological nuances and performance characteristics unique to anomaly synthesis algorithms. The absence of systematic evaluation frameworks creats critical gaps in understanding algorithmic strengths, operational boundaries and domain-specific adaptability-limitations. To address this critical gap, our reasearch focuses specifically on anomaly synthesis algorithms, with the primary objectives of establishing a systematic evaluation framework and conducting rigorous performance analysis. The proposed framework aims to provide scientific foundations for anomaly detection tasks, offering theoretical guidance for the selection and optimization of synthesis methods in practical applications.
@@
    There has been no systematic evaluation of different anomaly synthesis methods across various datasets, thereby lacking guidance for method selection in practical applications.
@@
    Data Ratio Impact: Investigate the impact of different sample proportions of synthsis anomalies on the performance of anomaly detection models.
@@
    Hybrid Strategies: Evaluate the effects of mixing abnormal data samples generated by mulitiple anomaly synthesis algorithms.
Figure 1: The overall workflow of ASBench. We disentangle the different stage in the anomaly synthesis and anomaly detection pipeline, and discuss the impact of variables at each stage on anomaly detection in separate subsections of Section IV.
As shown in Table 1, compared to other benchmark studies in the field of anomaly detection, our ASBench stands out for its unique features in anomaly synthesis decoupling, adjustable anomaly sample ratios, and diverse pipeline configurations. The overall workflow of ASBench, along with the chapter structure detailing analyses from various perspectives, is illustrated in Fig. 1.
@@ 2.2 Anomaly Synthesis
The development of anomaly generation techniques has progressed from simple image-layer operations to complex generation-based approaches.
Hand-crafted method is the mainstream research direction in the early stages of anomaly synthesis research. Initial methods primarily operated at the patch level for augmentation. Techniques such as CutOut [6] and RIAD [35] simulated anomalies by blacking out certain patches, while methods like CutPaste [17] further replaced patches with textures from other regions to mimic anomalies. However, the shapes of anomalies generated through these approaches often differed significantly from real anomalous shapes. Consequently, subsequent research gradually shifted toward using random noise to simulate anomalous regions. Representative methods include DRAEM [34], which employs Perlin noise [11] and external texture datasets to generate anomalous textures, and FractalAD [31], which utilizes fractal noise. Building upon these approaches, MemSeg [33] restricted anomalies to appear only in foreground regions, while NSA [25] used the object’s own texture to fill anomalous areas.
Generative-model-based methods have demonstrated outstanding performance, as the synthesized textures more closely resemble real anomalies and exhibit greater diversity. Early methods such as DFMGAN [9] and AdaBLDM [18] utilized GANs [24] and latent diffusion models to generate defect images, enhancing anomaly generation under small-sample conditions. The introduction of diffusion models [12], such as AnomalyDiffusion [13], enabled precise control over the spatial and visual embedding of anomalies, generating realistic anomalous images. RealNet [40] combined adaptive diffusion models with feature selection modules to enhance anomaly detection performance.
Based on the distinct characteristics of existing anomaly synthesis methods, we have selected patch augmentation-based and random noise-based methods from the manual synthesis category, as well as GANs-based and diffusion-based methods from the generation-model-based category for our benchmark.
@@ 3.1 Problem Definition
where $A$ represents the anomaly detection algorithm, $B$ denotes the anomaly synthesis method, $C$ corresponds to the industrial data set and $D$ specifies the proportion of synthetic abnormal samples in the training data. The $\otimes$ operator represents composability between components. This formulation establishes a systematic approach to evaluate the interdependencies between critical components.
@@ 3.2 Implementation Details
Random Shape Noise DestSeg [41], DRAEM [34], FractalAD [31], Memseg [33], NSA [25],
@@
Regarding evaluation metrics, we employ Area Under the Receiver Operating Characteristics (AU-ROC/AUC), Area Under Precision-Recall (AUPR/AP), and PRO to evaluate the abilities of anomaly localization. These five metrics collectively address both anomaly detection and segmentation performance at image and pixel granularities, ensuring comprehensive and precise evaluation.
@@ 4.1 Overall Comparison
Metric
AUC
Image
AP
Pixel
AUC
Image
AP
Pixel
AUC
Image
AP
Pixel
AUC
Image
AP
Pixel
AUC
Image
AP
Pixel
@@ 4.1.1 Comparative Experimental Framework
We conducted a comprehensive comparison across 5 datasets, 12 synthesis methods, and 4 detection algorithms, and the detailed results is presented in Table 4. To visualize the overall performance, we generated radar charts (Fig. 2) depicting AUC Image and AP Pixel aggregated from all datasets. Notably, for the overall evaluation, we adopted a weighted averaging approach using the number of subclasses within each dataset as weighting coefficients to account for inter-dataset categorical imbalances.
@@ 4.1.2 Analysis of Results
As evidenced by Fig. 3 and Table 4, significant performance variations are observed across different combinations of detection algorithm and synthesis method on various datasets. Regarding image-level anomaly classification, the four synthesis methods, DestSeg, DRAEM, Fractal, and AnomalyDiffusion exhibit generally stable performance across all four detection algorithms, with notable exceptions: CutPaste combined with DRAEM’s algorithm achieves state-of-the-art classification results on BTAD, while the integration of DestSeg synthesis and DRAEM detection proves optimal for VisA. These dataset-specific performance patterns underscore the importance of context-sensitive method selection in anomaly detection tasks. For pixel-level anomaly localization, the AnomalyDiffusion framework demonstrates robust compatibility with most synthesis methods on VisA, MPDD, BTAD, and MTD datasets. However, on the MVTec AD dataset, synthesis strategies derived from DRAEM and DestSeg yield superior localization performance compared to AnomalyDiffusion.
@@ 4.2.1 Comparative Experimental Framework
Figure 6: Comparison of detection pipelines: (a) DestSeg, (b) DRAEM, (c) MemSeg, and (d) Anomaly Detection.
Then, to analysis model robustness, we calculate the mean and variance of performance metrics for each synthesis method across different datasets and under different detection models. And we construct Fig. 6 to quantify the discrepancy in robustness among the methods.
@@ 4.2.2 Analysis of Results
As shown in the Fig. 2 and Fig. 5, different combinations of detection pipelines and synthesis methods exhibit significant performance variations across datasets. For anomaly localization, AnomalyDiffusion combined with most synthesis methods achieves strong performance on VisA, MPDD, BTAD, and MTD datasets; however, on the MVTec AD dataset, synthesis methods using DRAEM and DestSeg outperform AnomalyDiffusion. For image-level anomaly classification, the CutPaste combined with DRAEM detection pipeline achieves state-of-the-art performance on the BTAD dataset, while the DestSeg strategy paired with DRAEM yields optimal results on the VisA dataset.
Through comparative analysis of synthesis methods, we observe that approaches capable of generating complex and diverse anomalous regions significantly outperform handcrafted designs or those relying on real anomaly samples, and the potential advantages of real anomalies remaining underutilized. Handcrafted methods like such as CutPaste, CutOut, and FPI demonstrate inferior performance. Approaches leveraging real anomalies, including DFMGAN, RealNet, and AnomalyDiffusion, also do not exhibit significant advantages. The more effective methods, DRAEM, DestSeg, Fractal, and NSA, employ complex and diverse masks to increase anomaly diversity and mitigate overfitting risks. Defining anomalous regions through Fractal patterns or Perlin Noise proves more effective than random cropping or simulating real anomalies, as these techniques introduce greater stochasticity that improves model generalization capabilities.
@@ 4.3.1 Comparative Experimental Framework
This configuration enables dual analytical perspectives: (a) averaging across four detection models to evaluate how data ratios affect performance under different synthesis strategies, and (b) averaging across synthesis methods to examine how anomaly ratios impact individual model performance. On the other hand, we investigate the relationship between synthesis method stability and performance by analyzing how the performance differential (between optimal and worst-case scenarios) varies with anomaly sample ratios. The model-averaged performance results and the performance differentials are recorded in Table 8 and 9. Based on tabulated mean values and performance gaps, we generate scatter plots and calculate the related correlation coefficients to analyse.
@@
MVTec 0.5850 0.7425 0.7968 0.7985
BTAD 0.5421 0.7578 0.6277 0.8028
VisA 0.5276 0.7462 0.6852 0.5666
MPDD 0.3735 0.7808 0.7832 0.7642
MTD 0.6345 0.7401 0.5620 0.4814
Table 6: Image AUROC of percent 0 for various datasets.
Dataset DestSeg DRAEM MemSeg AnomalyDiffusion
MVTec 0.0362 0.0531 0.1440 0.0730
BTAD 0.0330 0.0572 0.0442 0.0361
VisA 0.0071 0.0046 0.0082 0.0500
MPDD 0.0345 0.0756 0.1411 0.1676
MTD 0.0820 0.0742 0.0743 0.0776
Table 7: Pixel AUPR of percent 0 for various datasets.
@@ 4.3.2 Analysis of Results
As shown in Table 6 and 7 along with Fig. 8, the experimental results under anomaly-free conditions demonstrate significantly inferior detection performance compared to scenarios containing anomalous samples. The significant difference observed demonstrate the efficacy of anomaly synthesis in enhancing detection capabilities across these algorithms. However, Fig. 10 illustrates that, after introducing synthetic anomalies, different ratios exhibit minimal performance variations, and no statistically significant optimal ratio point has been identified through current comparative analyses.
@@ 4.4.1 Comparative Experimental Framework
The analytical framework employs two-dimensional scatter plots, Fig. 11 visualizing relationships between generation metrics and detection performance indicators, accompanied by Pearson correlation coefficients with statistical significance levels (p-values), as detailed in Table 10 and 11.
@@
Dataset MVTec-AD BTAD VisA MPDD MTD
SSIM vs AUPR -0.1280 0.6918 0.2360 0.4602 0.0243 0.4602 -0.0236 0.9420 0.2330 0.4662
PSNR vs AUPR -0.2662 0.4030 0.2079 0.5168 -0.1908 0.5168 -0.2388 0.4548 0.1032 0.7495
LPIPS vs AUPR 0.2171 0.4978 -0.2658 0.4038 0.3083 0.4038 -0.0520 0.8725 -0.2515 0.4305
IS vs AUPR 0.5494 0.0643 -0.0120 0.9705 0.5687 0.9705 0.2882 0.3636 0.0086 0.9787
FID vs AUPR 0.2373 0.4576 -0.5014 0.0968 0.2724 0.0968 0.1310 0.6848 -0.1109 0.7316
KID vs AUPR -0.1693 0.5988 -0.4766 0.1172 -0.0043 0.1172 -0.0136 0.9666 -0.1620 0.6149
Table 11: Correlation and P-values between six metrics and pixel-level AUPR across datasets.

### v3

@@ Abstract
Anomaly detection plays a pivotal role in manufacturing quality control, yet its application is constrained by limited abnormal samples and high manual annotation costs. While anomaly synthesis offers a promising solution, existing studies predominantly treat anomaly synthesis as an auxiliary component within anomaly detection frameworks, lacking systematic evaluation of anomaly synthesis algorithms. Current research also overlooks crucial factors specific to anomaly synthesis, such as decoupling its impact from detection, quantitative analysis of synthetic data and adaptability across different scenarios. To address these limitations, we propose ASBench, the first comprehensive benchmarking framework dedicated to evaluating anomaly synthesis methods. Our framework introduces four critical evaluation dimensions: (i) the generalization performance across different datasets and pipelines (ii) the ratio of synthetic to real data (iii) the correlation between intrinsic metrics of synthesis images and anomaly detection performance metrics, and (iv) strategies for hybrid anomaly synthesis methods. Through extensive experiments, ASBench not only reveals limitations in current anomaly synthesis methods but also provides actionable insights for future research directions in anomaly synthesis. Code is available at https://github.com/M-3LAB/ASBench.
{IEEEImpStatement}
Industrial image anomaly detection is a highly popular field in both academia and industry. However, academic research on industrial image anomaly detection primarily focuses on unsupervised anomaly detection, whereas industrial applications often directly employ supervised training methods to develop the required models. Anomaly synthesis can effectively bridge the gap between academia and industry by transforming unsupervised anomaly detection into supervised model training. Our paper presents the first comprehensive benchmark for analyzing anomaly synthesis in industrial images. Through a detailed examination of various anomaly synthesis methods, we aim to identify the currently optimal synthesis method and narrow the gap between industrial practices and academic research.
@@
Anomaly detection, Defect detection, Unsupervised learning
@@ 1 Introduction
Anomaly detection has emerged as a critical technology in modern manufacturing quality control and healthcare monitoring, playing an important role in ensuring product reliability and enhancing diagnostic accuracy [7] [32]. Despite its significance, the practical deployment of anomaly detection systems is often constrained by two major challenges: the scarcity of abnormal samples and the prohibitively high costs associated with manual annotation. These limitations significantly impede the performance and scalability of the detection models. To address these issues, anomaly synthesis has gained traction as a promising solution, offering the capability to generate representative abnormal samples. This approach not only alleviates the data scarcity problem but also substantially reduces the reliance on costly manual annotation efforts. Jan and Christian (2022) [9] summarize the anomaly synthesis methods. The significant potential has been demonstrated in various domains, including surface defect detection and medical image analysis. Anomaly synthesis can be effectively integrated with real anomaly data or directly applied to enhance detection systems.
While anomaly synthesis holds transformative potential for advancing anomaly detection capabilities, current anomaly synthesis methods remain largely confinded to auxiliary roles within anomaly detection frameworks. This paradigm neglects rigorous benchmarking and detailed exploration of the methodological nuances and performance characteristics unique to anomaly synthesis algorithms. The absence of systematic evaluation frameworks creates critical gaps in understanding algorithmic strengths, operational boundaries and domain-specific adaptability-limitations. To address this critical gap, our reasearch focuses specifically on anomaly synthesis algorithms, with the primary objectives of establishing a systematic evaluation framework and conducting rigorous performance analysis. The proposed framework aims to provide scientific foundations for anomaly detection tasks, offering theoretical guidance for the selection and optimization of synthesis methods in practical applications.
@@
    There has been no systematic evaluation of different anomaly synthesis methods across various datasets, thereby lacking guidance in practical applications.
@@
    Data Ratio Impact: Investigate the impact of different sample proportions of synthesis anomalies on the performance of anomaly detection models.
@@
    Hybrid Strategies: Evaluate the effects of mixing abnormal data samples generated by multiple anomaly synthesis algorithms.
Figure 1: The overall workflow of ASBench. We disentangle different stages in the anomaly synthesis and anomaly detection pipeline, and discuss the impact of variables at each stage on anomaly detection in separate subsections of Section IV.
As shown in Table 1, compared to other benchmark studies in the field of anomaly detection, our ASBench stands out for its unique features in anomaly synthesis decoupling, adjustable anomaly sample ratios, and diverse pipeline configurations. The overall workflow of ASBench, along with the chapter structure detailing analyzes from various perspectives, is illustrated in Fig. 1.
@@
The remainder of this paper is structured as follows. Section II reviews related literature on anomaly detection pipelines and anomaly synthesis methods. Section III introduces the experimental settings of our proposed benchmark, ASBench. In Section IV, a comprehensive experimental evaluation is presented along with a discussion of the results from four perspectives. Finally, Section V concludes the paper.
@@ 2.1 Anomaly Detection
The field of anomaly detection has seen excellent development in available datasets. MVTec AD [2], as the first comprehensive dataset, covering multiple object and defect types, has become the widely recognized benchmark in the field. MTD [16] is specifically designed for saliency detection in tile images, focusing on the identification of anomalies at a micro-scale. BTAD [25] is a dataset from real industrial scenarios where images have relatively simple backgrounds, making the detection task more straightforward. MPPD [17] focuses on defects in metal painted parts during the manufacturing process, emphasizing detection capability under diverse viewing angles and complex conditions. The innovation of the VisA dataset [47] lies in its introduction of a multi-instance sample setting. These datasets align with the objectives of the anomaly detection tasks targeted by ASBench from various perspectives.
PAD [45] and Real-IAD [33] share a key characteristic with MPPD by focusing on multi-viewpoint imaging to closely approximate real industrial conditions. Comprehensive 3D datasets such as MVTec 3D-AD [4] and Real3D-AD [23] are primarily oriented towards anomaly detection and localization in 3D data, which does not fully align with ASBench’s current focus on 2D image detection tasks.
@@ 2.2 Anomaly Synthesis
The development of anomaly generation techniques has evolved from simple image‑level manipulations to complex generative approaches. Hand-crafted method is the mainstream research direction in the early stages of anomaly synthesis research. Initial methods primarily operated at the patch level for augmentation. Techniques such as CutOut [8] and RIAD [38] simulated anomalies by blacking out certain patches, while methods like CutPaste [19] further replaced patches with textures from other regions to mimic anomalies. However, the shapes of anomalies generated through these approaches often differed significantly from real anomalous shapes. Consequently, subsequent research gradually shifted toward using random noise to simulate anomalous regions. Representative methods include DRAEM [37], which employs Perlin noise [13] and external texture datasets to generate anomalous textures, and FractalAD [34], which utilizes fractal noise. Building upon these approaches, MemSeg[36] constrained anomalies to appear only in foreground regions, whereas NSA [28] filled anomalous areas with the object’s own texture via Poisson image editing.
With the advancement of generative models, methods based on generative models have gradually become mainstream in anomaly synthesis, as they produce textures that are closer to real anomalies and offer greater diversity. Early methods such as DFMGAN [11] and AdaBLDM [20] utilized GANs [27] and latent diffusion models to generate defective images, thereby improving anomaly generation capabilities under few‑shot conditions. In recent years, diffusion models [14] have been introduced to the anomaly synthesis task. Methods such as AnomalyDiffusion [15] enable precise control over the spatial distribution and visual appearance of anomalies, capable of generating more realistic anomalous images. RealNet [43] combined adaptive diffusion models with feature selection modules to enhance anomaly detection performance.
Based on the distinct characteristics of existing anomaly synthesis methods, we have selected patch augmentation-based and random noise-based methods from the Hand-crafted synthesis category, as well as GANs-based and diffusion-based methods from the generation-model-based category for our benchmark.
@@ 3.1 Problem Definition
where AA represents the anomaly detection algorithm, $B$ denotes the anomaly synthesis method, $C$ corresponds to the industrial dataset and $D$ specifies the proportion of synthetic abnormal samples in the training data. The $\otimes$ operator represents composability between components. This formulation establishes a systematic approach to evaluate the interdependencies between critical components.
@@ 3.2 Implementation Details
Random Shape Noise DestSeg [44], DRAEM [37], FractalAD [34], MemSeg [36], NSA [28]
@@
Regarding evaluation metrics, we employ Area Under the Receiver Operating Characteristics (AU-ROC/AUC), Area Under Precision-Recall (AUPR/AP), and Per-Region Overlap (PRO) [1] to evaluate the abilities of anomaly localization. These five metrics collectively address both anomaly detection and segmentation performance at image and pixel granularities, ensuring comprehensive and precise evaluation.
@@ 4.1 Overall Comparison
Metric AUC Image AP Pixel AUC Image AP Pixel AUC Image AP Pixel AUC Image AP Pixel AUC Image AP Pixel
@@ 4.1.1 Comparative Experimental Framework
We conducted a comprehensive comparison across 5 datasets, 12 synthesis methods, and 4 detection algorithms, and the detailed results are presented in Table 4. To visualize the overall performance, we generated radar charts (Fig. 2) depicting AUC Image and AP Pixel aggregated from all datasets. Notably, for the overall evaluation, we adopted a weighted averaging approach using the number of subclasses within each dataset as weighting coefficients to account for inter-dataset categorical imbalances.
@@ 4.1.2 Analysis of Results
As evidenced by Fig. 3 and Table 4, significant performance variations are observed across different combinations of detection algorithm and synthesis method on various datasets. Regarding image-level anomaly classification, the four synthesis methods, DestSeg, DRAEM, Fractal, and AnomalyDiffusion exhibit generally stable performance across all four detection algorithms. With notable exceptions: CutPaste combined with DRAEM’s algorithm achieves state-of-the-art results on BTAD, where textures are relatively simple and homogeneous. In contrast, the VisA dataset contains multi-instance objects, making it more suitable for methods like FPI and NSA that employ seamless image editing and blending, or Fractal-based techniques that generate diversified anomalies through synthetic fractal patterns. For datasets focusing on multi-view, fine-grained details (e.g., MPDD, MTD), generative models such as AnomalyDiffusion and RealNet perform better, as they excel at simulating detailed anomalies. For pixel-level anomaly localization, the AnomalyDiffusion detection pipelines demonstrates robust compatibility with most synthesis methods on VisA, MPDD, BTAD, and MTD datasets. However, on the MVTec AD dataset, synthesis strategies derived from DRAEM and DestSeg yield superior localization performance compared to AnomalyDiffusion.
@@ 4.2.1 Comparative Experimental Framework
Figure 6: Comparison of detection pipelines: (a) DestSeg, (b) DRAEM, (c) MemSeg, and (d) Anomaly Diffusion.
Then, to analyze model robustness, we calculate the mean and variance of performance metrics for each synthesis method across different datasets and under different detection models. And we construct Fig. 6 to quantify the discrepancy in robustness among the methods.
@@ 4.2.2 Analysis of Results
As shown in the Fig. 2 and Fig. 5, through comparative analysis of synthesis methods, we observe that approaches capable of generating complex and diverse anomalous regions significantly outperform handcrafted designs or those relying on real anomaly samples, and the potential advantages of real anomalies remaining underutilized. Handcrafted methods like such as CutPaste, CutOut, and FPI demonstrate inferior performance. Approaches leveraging real anomalies, including DFMGAN, RealNet, and AnomalyDiffusion, also do not exhibit significant advantages. The more effective methods, DRAEM, DestSeg, Fractal, and NSA, employ complex and diverse masks to increase anomaly diversity and mitigate overfitting risks. Defining anomalous regions through Fractal patterns or Perlin Noise proves more effective than random cropping or simulating real anomalies, as these techniques introduce greater stochasticity that improves model generalization capabilities.
@@ 4.3.1 Comparative Experimental Framework
This configuration enables dual analytical perspectives: (a) averaging across four detection models to evaluate how data ratios affect performance under different synthesis strategies, and (b) averaging across synthesis methods to examine how anomaly ratios impact individual model performance. On the other hand, we investigate the relationship between synthesis method stability and performance by analyzing how the performance differential (between optimal and worst-case scenarios) varies with anomaly sample ratios. The model-averaged performance results and the performance differentials are recorded in Table 7 and 8. Based on tabulated mean values and performance gaps, we generate scatter plots and calculate the related correlation coefficients to analyze.
@@
Metric AUROC / AUPR / PRO AUROC / AUPR / PRO AUROC / AUPR / PRO AUROC / AUPR / PRO
MVTec AD 0.5850 / 0.0362 / 0.1536 0.7425 / 0.0531 / 0.2324 0.7968 / 0.1440 / 0.5065 0.7985 / 0.0730 / 0.3416
BTAD 0.5421 / 0.0330 / 0.1572 0.7578 / 0.0572 / 0.2436 0.6277 / 0.0442 / 0.4859 0.8028 / 0.0361 / 0.2822
VisA 0.5276 / 0.0071 / 0.1850 0.7462 / 0.0046 / 0.3265 0.6852 / 0.0082 / 0.4862 0.5666 / 0.0500 / 0.6029
MPDD 0.3735 / 0.0345 / 0.1874 0.7808 / 0.0756 / 0.2718 0.7832 / 0.1411 / 0.4326 0.7642 / 0.1676 / 0.6418
MTD 0.6345 / 0.0820 / 0.1646 0.7401 / 0.0742 / 0.2150 0.5620 / 0.0743 / 0.4753 0.4814 / 0.0776 / 0.2921
Table 6: Performance comparison (AUROC/AUPR/PRO) of various methods under anomaly-free conditions on different datasets.
@@ 4.3.2 Analysis of Results
As shown in Table 6 and Fig. 8, the experimental results under anomaly-free conditions demonstrate significantly inferior detection performance compared to scenarios containing anomalous samples. The significant difference observed demonstrate the efficacy of anomaly synthesis in enhancing detection capabilities across these algorithms. However, Fig. 10 illustrates that, after introducing synthetic anomalies, different ratios exhibit minimal performance variations, and no statistically significant optimal ratio point has been identified through current comparative analyzes.
@@ 4.4.1 Comparative Experimental Framework
The analytical framework employs two-dimensional scatter plots, Fig. 11 visualizing relationships between generation metrics and detection performance indicators, accompanied by Pearson correlation coefficients with statistical significance levels (p-values), as detailed in Table 9 .